# SWAN: Preprocessing SGD Enables Adam-Level Performance On LLM Training With Significant Memory Reduction

### v1

@@ 5.3 SWAN Performance on LLM Pre-training Tasks
Finally, we present results on the task of pre-training large language models. Below we consider LLaMa architecture used in Zhao et al. (2024a) with RMSNorm and SwiGLU activations. We consider sizes of 60M, 130M, 350M and 1.3B parameters on C4 dataset Raffel et al. (2020), trained with an effective batch size of 130K tokens (512 batchsize and 256 context length). SWAN is applied to all linear modules (in both attention and MLP layers). We run all experiments with BF16 format by default.
@@
Since the main goal of SWAN is to explore the feasibility of EMA-free adaptive optimizers for LLMs, we mainly compare to Adam and its memory efficient variants. We use the exact evaluation settings of Zhao et al. (2024a); along with Adam, these include:
  • Galore Zhao et al. (2024a): the state-of-the-art memory efficient variant of Adam that based on low-rank projections of gradients.
  • Low-Rank (Kamalakara et al., 2022): Low-rank reparameterization of the weights with the trainable factorization ${\mathbf{W}}={\mathbf{B}}{\mathbf{A}}$
  • LoRA: LoRA Hu et al. (2021) applied for pre-training in the exact setting as used in Zhao et al. (2024a).
  • ReLoRA Lialin et al. (2023): a full-rank learning extension of Lora with parameter merging.
@@
For all baselines, we directly apply the same hyperparameter configs (including model configs, optimizers, schedulers) open sourced by Zhao et al. (2024a)  [^4] This does not include Adam, since the authors of Zhao et al. (2024a) did not disclose their Adam configs. However, we used the same learning rate sweep procedure as described in their paper. This result in an optimal learning rate of 0.001 for 60M, 130M and 350M models, and 0.0007 for 1.3 B models. These are also consistent with the optimal learning rates found in some other works, such as Pagliardini et al. (2024). . This includes method-specific fine-tuned optimal learning rates obtained via searching on fixed grid Zhao et al. (2024a). For our SWAN method, we simply use a hand-picked default hyperparameter setting out-of-the-box, without fine tuning. These include a un-tuned, default learning rate of 0.001 across all models, and the 10 steps of Newton-Schultz iteration. Notably, thanks to the strong robustness offered by $\mathtt{GradWhitening}$ to ill-conditioned problems, SWAN can safely start training without any learning rate warm-up. This is not the case for other baselines such as Adam, in which they failed to converge without long warm-ups in our ablation experiments. Apart from the warm-up hyperparameter, all the rest of the parameters in learning rate scheduling used by SWAN is the same as all other baselines.
Remark on fair comparison
As empirically shown in Kaddour et al. (2024), it is essential to use the same training budget when comparing efficient optimizers; unfortunately, this fact was not widely respected in some recent work of efficient algorithms. Kaddour et al. (2024) showed that certain state-of-the-art efficient optimizers has shown no gain over Adam under the same training budget. Therefore, in our experiments above, we exactly follow the rule of equal training budget (in our case, defined in terms number of tokens), and forces all methods to have learning rate fully decayed to the same targe learning rate at the end of optimization. For a even fairer comparison, we left the learning rates of SWAN untuned.
@@ Results
Results are shown in Table 1 and Figure 6. SWAN outperforms all baselines in both validation PPLs, as well as estimated memory consumption (SWAN theoretically should consume the same memory as SGD). Furthermore, to our surprise, on both 350M and 1.3B models, SWAN achieved $>$ 2X speed up against Adam in terms of number of steps (hence number of tokens), see Figure 1 and Figure 7 (a). More in-depth discussion of speed-up will be provided in Section 5.4.
@@ 5.4 Is the Improvement Multiplicative or Additive?
One natural question when evaluating speedup factors is whether our method provides a multiplicative or additive speedup over Adam. Specifically, multiplicative speedup indicates an optimizer’s performance advantage over a baseline increases proportionally with time, maintaining a consistent speed-up ratio. This is the ideal scenario we would like to achieve. In contrast, additive speed-up implies only a fixed, constant advantage in steps over baseline, which is less desired. To investigate this, we utilize two plots: the speed-up ratio comparison and the perplexity comparison (Figure 7), across different model sizes. Before understanding those plots, we first introduce the following concepts:
@@
The speedup ratio $R(P)$ is defined as the ratio of the number of training steps Adam requires to reach a specific evaluation perplexity (PPL) to the number of steps SWAN requires to achieve the same PPL. Mathematically, for a given PPL threshold $P$, the speedup ratio is expressed as:
$R(P)=\frac{S_{\text{Adam}}(P)}{S_{\text{SWAN{}}}(P)}$ (6)
where $S_{\text{Adam}}(P)$ and $S_{\text{SWAN{}}}(P)$ denote the training steps taken by Adam and SWAN, respectively, to reach perplexity $P$.
@@
To determine if the speed-up is additive, we estimate a counterfactual additive curve by assuming a constant step advantage $\Delta$ of SWAN over Adam. This is calculated using the first 10% - 20% (depending on the number of maximum steps) of the training data:
$\Delta=\frac{1}{N}\sum_{i=1}^{N}\left(S_{\text{Adam}}(P_{i})-S_{\text{SWAN{}}}(P_{i})\right)$ (7)
where $N$ is the number of PPL thresholds within the initial 10% - 20% of training steps.
For the speedup ratio plot:
$R_{\text{additive}}(P)=\frac{S_{\text{Adam}}(P)}{S_{\text{Adam}}(P)-\Delta}$ (8)
For the perplexity comparison plot:
$\text{PPL}_{\text{additive}}(S)=\text{PPL}_{\text{Adam}}(S+\Delta)$ (9)
@@ Results
Results are shown in Figure 7. In both subplots, the actual performance of SWAN is compared against the counterfactual additive curves. If the actual plot lies significantly above the counterfactual additive curve, it indicates a stronger tendency towards multiplicative speedup, meaning SWAN’s performance gains are more than just a fixed step advantage. Analyzing models of sizes 60M, 130M, 350M, and 1.3B parameters, we observe that:
  • For smaller models (60M and 130M parameters), the actual speedup ratios closely follow the counterfactual additive curves, suggesting that the speedup of SWAN over Adam is primarily additive.
  • For larger models (350M and 1.3B parameters), the actual plots significantly exceed the counterfactual additive curves, and has a clear sign of accelerated growth. This pronounced deviation indicates a multiplicative speedup, demonstrating that SWAN offers increasingly substantial efficiency gains as model complexity scales.
@@ 5.5 Memory Efficiency and Throughput Analysis
This section, we benchmark the memory footprint of SWAN as well as its throughput.
@@
We benchmark the memory footprint of SWAN, Adam, and Galore in a practical scenario.. Unlike Zhao et al. (2024a) where memory is benchmarked under a layer-wise training strategy, we directly measure end-to-end memory under full-model training, with batch size = 1 sequence, across different model sizes (1.3B, 7B, and 13B). Results are shown in Figure 1 (c), SWAN achieves near-theoretical limit of optimizer memory efficiency of full-parameter training (assuming no quantization). This means that SWAN can reproduce the memory footprint of vanilla SGD, reaching $\approx 50\%$ reduction on total memory, and $\approx 100\%$ reduction on optimizer states. This result again highlights the advantage of the stateless design.
@@
To measure the throughput of SWAN, we consider training a 1.3 B model on 4 $\times$ A100, under constant batch size = 130K tokens. Here We present two metrics: absolute throughput, measured by number of training tokens consumed per second; and effective throughput, which is absolute throughput adjusted by the efficiency of optimizer relative to Adam. The purpose of the first metric (absolute throughput) is to measure whether SWAN (especially the use of N-S iterations in $\mathtt{GradWhitening}$) create a significant drop in training throughput; and the second metric (effective throughput) takes into account the fact that different optimizers utilizes training tokens at different efficiency level. Results are shown in Figure 1 (d), SWAN closely reproduces the absolute throughput of Adam, implying that unlike full SVD used in (Zhao et al., 2024a), the computational overhead of $\mathtt{GradWhitening}$ is minimal. As a result, SWAN has doubled the effective throughput vs Adam, due to its high token utilization efficiency as shown in Section 5.3. All results are obtained using $\mathtt{GradWhitening}$ in BF16 format.
@@ 5.6 Ablation Studies
How does $\mathtt{GradNorm}$ and $\mathtt{GradWhitening}$ contribute to the performance boost?
We consider the following 6 ablation settings: SWAN(full), SWAN ($\mathtt{GradNorm}$ only), SWAN ($\mathtt{GradWhitening}$ only), Adam (full), Adam (momentum only), and Adam (second moment only). Results are shown in Figure 8 (a). It suggests that both $\mathtt{GradNorm}$ and $\mathtt{GradWhitening}$ makes significant contribution towards the final performance of SWAN; and removing any of them results in suboptimal performance. This is similar to how the first and second moment EMAs of Adam jointly contributes to the final performance.
@@
Is it true that SWAN works only because its $\mathtt{GradNorm}$ component might have increased the effective learning rates?
To answer this question, we remove the $\mathtt{GradNorm}$ component of SWAN, and perform learning rate sweep for it. We start from the default base learning rate used by full SWAN, and then scale it with multipliers ranging from $1$ to $10^{3}$. Results are shown in Figure 8 (b), where the final validation ppl of SWAN without $\mathtt{GradNorm}$ under different learning rate multipliers are shown in blue line; and the performance of the full SWAN under default learning rate is shown as a dashed red reference line. The result suggests that, although increasing the effective learning rates might improve the performance of SWAN without $\mathtt{GradNorm}$, there is still a significant cap between between SWAN without $\mathtt{GradNorm}$ and the full SWAN. Together with results in Section 5.1, we can conclude that the gradient noise stabilization of $\mathtt{GradNorm}$ is indeed essential, and it cannot be trivialized as larger effective learning rates.
@@
In Section 5.3, we showed that SWAN does not need any learning rate warm-ups and can train with large learning rates (the default 0.001 learning rate used by all methods is considerably large in LLMs settings). Here we perform ablation and compare Adam and SWAN both under warm-ups and not warm-ups. As suggested by the results in Figure 8 (c), we found that for SWAN, training without learning rate warm-ups indeed gives better performance. Meanwhile, SWAN still outperforms Adam (with warm-up) under the same warm-up schedule of Adam. On the contrary, Adam only works with proper learning rate warmup. If we remove warmups, the performance of Adam becomes drastically worse. This confirms our claim that SWAN is more robust to the local geometries of the optimization problems and can work without (and with) learning rate warm ups.
6 Related works
Low-rank memory efficient methods
Low-rank memory-efficient optimizers have become pivotal in training large language models (LLMs) by reducing memory consumption while maintaining model performance. A foundational approach in this domain is LORA introduced by Hu et al. (2021), which fine-tunes pre-trained models using low-rank adaptors for each layer, thereby minimizing the memory footprint. This leads a full line of low-rank adaptation research, but they are mostly focused on post-training adaptations. More recently, ReLoRA Lialin et al. (2023),FLORA Hao et al. (2024) and Galore Zhao et al. (2024a) further push the boundaries of memory-efficient pre-training for LLMs, by introducing the idea of using low-rank projection of gradients to achieve full-rank learning. They have demonstrated significant memory savings on optimizer states, with relatively low impact on model performance. However, these methods often require 30–50 % of the full Adam’s states and may underperform in terms of convergence speed and final model quality (Zhao et al., 2024a). Moreover, it is still an open question on the design choices of these methods, for example: how and why to choose low-rank subspace and projectors? Are these projections properly used, given the non-exchangeability nature between low-rank projection and Adam EMA operator? Why spikes in loss curves are observed during subspace are shifted/mixed (Chen et al., 2024)? These unaddressed questions might make the behavior of low-rank gradient optimizers unpredictable in practice.
Non-adaptive methods with partial EMA removal.
Lion (Chen et al., 2023) and Signum (Bernstein et al., 2018) have introduced effective ways to simplify optimization by partially removing the exponential moving average (EMA) of the second moment. These methods reduce memory and computational overhead while retaining the benefits of momentum through the first moment EMA, offering a practical trade-off between simplicity and performance. However, their reliance on momentum inherently limits their extension to further remove both first and second moment EMAs are removed, as both would reduce to SignSGD (Bernstein et al., 2018), which lacks the adaptive learning rate capability of Adam. This might explain why sign-based methods have limited success in LLM pre-training tasks (Chen et al., 2023; Liu et al., 2023). On the contrary, our approach completely removes all EMA terms while maintaining the adaptive learning rate nature, which is the key to the success of modern adaptive optimizers.

### v2

@@ 6.1 SWAN Performance on LLM Pre-training Tasks
We evaluate SWAN on large language model (LLM) pre-training tasks using a LLaMa-based architecture Touvron et al. (2023a) with RMSNorm and SwiGLU activations (Zhang & Sennrich, 2019; Gao et al., 2023). We consider models with 60M, 130M, 350M, and 1.3B parameters, all trained on the C4 dataset Raffel et al. (2020) using an effective token batch size of 130K tokens (batch size 512, context length 256). Following the setup of Zhao et al. (2024a), SWAN is applied to all linear modules in both attention and MLP blocks. Training uses BF16 precision by default. The other evaluation settings are the same as Zhao et al. (2024a).
@@
Since SWAN aims to explore EMA-free adaptive optimizers for LLMs, we focus on comparing it with Adam and other low-rank optimizers. The baselines are:
  • Adam (Kingma & Ba, 2015): one standard choice for training large models.
  • Galore Zhao et al. (2024a): a memory-efficient Adam variant with low-rank gradient projections.
  • Low-Rank (Kamalakara et al., 2022): a low-rank factorization approach (${\mathbf{W}}={\mathbf{B}}{\mathbf{A}}$).
  • LoRA Hu et al. (2021): the LoRA method applied for pre-training as in Zhao et al. (2024a).
  • ReLoRA Lialin et al. (2023): a full-rank extension of LoRA with parameter merging.
@@
We use the publicly available hyperparameter configurations from Zhao et al. (2024a) for all baselines [^2] This does not include Adam, since the authors of Zhao et al. (2024a) did not disclose their Adam configs. However, we used the same learning rate sweep procedure as described in their paper. This result in an optimal learning rate of 0.001 for 60M, 130M and 350M models, and 0.0007 for 1.3 B models., including method-specific learning rates selected through a grid search. For SWAN, we use a default setting without extensive tuning. We match the global learning rate (0.001) used by Adam, except for the 1.3B model (Adam diverges at 0.001, see Appendix H). We use 10 steps of Newton–Schulz iteration for $\mathtt{GradWhitening}$. Unlike Adam, SWAN does not require a warm-up phase to ensure stable training. All other scheduling parameters remain the same as in the baselines. Full details can be found in Appendix H.
Fair Comparison
We follow Kaddour et al. (2024) by allocating the same total training budget (in terms of processed tokens) for all methods. By the end of training, all methods decay to the same target learning rate. We also leave SWAN’s learning rates largely untuned for a more direct comparison.
@@ Results
Table 1 and Figure 4 show the results. SWAN achieves lower validation perplexities than the baselines with lower estimated memory consumptions. On the 350M and 1.3B models, SWAN reaches over a 2$\times$ speedup (in steps or tokens) relative to Adam (Figure 1 and Figure 5 (a)). Further details are in Section 6.2.
@@ 6.2 Is the Improvement Multiplicative or Additive?
A key question in assessing speedup factors is whether the improvement over Adam is multiplicative or additive. A multiplicative speedup implies that the optimizer’s relative advantage remains proportionally consistent over time, while an additive speedup suggests a less desired constant step advantage. We explore these scenarios using two plots, the speed-up ratio comparison and the perplexity comparison (Figure 5), across different model sizes.
@@
We define the speedup ratio $R(P)$ for a given perplexity (PPL) threshold $P$ as the ratio of the number of training steps Adam requires to reach a specific evaluation perplexity (PPL) to the number of required steps for SWAN:
$R(P)=\frac{S_{\text{Adam}}(P)}{S_{\text{SWAN{}}}(P)},$ (6)
where $S_{\text{Adam}}(P)$ and $S_{\text{SWAN{}}}(P)$ are the training steps required by Adam and SWAN, respectively, to reach perplexity $P$.
@@
To test whether the speedup is additive, we compute a counterfactual additive curve by assuming SWAN gains a fixed step advantage $\Delta$ over Adam in early training (approximately the first 10%–20% of total steps):
$\Delta=\frac{1}{N}\sum_{i=1}^{N}\left(S_{\text{Adam}}(P_{i})-S_{\text{SWAN{}}}(P_{i})\right),$ (7)
where $N$ is the number of PPL thresholds considered. We then use $\Delta$ to define the counterfactual additive speedup ratio:
$R_{\text{additive}}(P)=\frac{S_{\text{Adam}}(P)}{S_{\text{Adam}}(P)-\Delta},$ (8)
and the counterfactual additive perplexity estimate:
$\text{PPL}_{\text{additive}}(S)=\text{PPL}_{\text{Adam}}\left(S+\Delta\right).$ (9)
@@ Results
Figure 5 compares SWAN’s actual performance against the counterfactual additive curves. If SWAN’s actual curves exceed these additive estimates, it indicates a tendency towards a multiplicative speedup, instead of the additive advantage. We summarize the observations for model sizes of 60M, 130M, 350M, and 1.3B parameters:
  • For smaller models (60M and 130M), the actual speedup trajectories align closely with the additive baseline, indicating a primarily additive speedup.
  • For larger models (350M and 1.3B), the actual curves rise noticeably above the additive estimates, suggesting a multiplicative speedup. This indicates that SWAN yields increasing efficiency gains as model size grows.
@@ 6.3 Memory Efficiency and Throughput Analysis
In this section, we measure both the memory footprint and throughput of SWAN.
@@
We compare SWAN, Adam, and Galore on a single A100 GPU. Unlike Zhao et al. (2024a), which reports layer-wise training memory usage, we measure total end-to-end memory consumption under full-model training using batch size of $1$ for LLaMA with 1.3B, 7B, and 13B parameters. As shown in Figure 1 (c), SWAN’s memory usage is close to that of vanilla SGD, providing nearly a 50% reduction in total memory by removing the needs for first and second moments compared to Adam. This observation underlines the benefit of a stateless design.
@@
We also assess throughput when training a 1.3B-parameter model on 4 A100 GPUs with a token batch size of 130K tokens. We use two metrics:
  • Absolute throughput: number of tokens processed per second.
  • Effective throughput: absolute throughput adjusted by SWAN’s token efficiency relative to Adam.
These metrics evaluate whether $\mathtt{GradWhitening}$ (which uses Newton-Schulz iterations) significantly reduces training speed and also account for the fact that some optimizers make more effective use of training tokens. As shown in Figure 1 (d), SWAN achieves an absolute throughput comparable to Adam, indicating that the computational overhead of $\mathtt{GradWhitening}$ is minimal. Consequently, SWAN exhibits a higher effective throughput than Adam because of its more efficient use of tokens during training. All experiments use BF16 precision for $\mathtt{GradWhitening}$.
@@ 6.4 Ablation Studies
Effect of $\mathtt{GradNorm}$ and $\mathtt{GradWhitening}$ on Performance
We consider six ablation settings: (1) SWAN(full), (2) SWAN ($\mathtt{GradNorm}$ only), (3) SWAN ($\mathtt{GradWhitening}$ only), (4) Adam (full), (5) Adam (momentum only), and (6) Adam (second moment only). As shown in Figure 6 (a), both $\mathtt{GradNorm}$ and $\mathtt{GradWhitening}$ contribute to SWAN’s final performance. Removing either results in performance degradation. Similarly, Adam also requires first and second moment estimates for optimal performance.
@@
Does SWAN Succeed Only by Increasing Effective Learning Rates though $\mathtt{GradNorm}$?
To answer this question, we remove the $\mathtt{GradNorm}$ operator from SWAN and run a learning rate sweep. Starting with the default learning rate of full SWAN, we apply multipliers from $1$ to $10^{3}$. In Figure 6 (b), the blue line shows the final validation perplexity for SWAN without $\mathtt{GradNorm}$ at different learning rates multipliers, while the dashed red line represents the performance of full SWAN at the default learning rate. Although raising the learning rate can improve the performance of SWAN without $\mathtt{GradNorm}$, the gap to full SWAN remains large. This indicates that $\mathtt{GradNorm}$’s gradient noise stabilization is essential and cannot be replaced simply by increasing the learning rate.
@@
Section 6.1 shows that SWAN can train with no warm-up phase even under a relatively large learning rate (0.001). Here, we compare Adam and SWAN with and without warm-ups. As seen in Figure 6 (c), SWAN without the warm-up phase gives better performance, and it still outperforms Adam under Adam’s own warm-up schedule. On the other hand, Adam’s performance decreases drastically without a proper warm-up. These findings suggest that SWAN is more robust to warm-up schedule and can train effectively with or without it.
6.5 Analysis of $\mathtt{GradNorm}$ and $\mathtt{GradWhitening}$
6.5.1 Does $\mathtt{GradNorm}$ Stabilize Gradient Distributions of SGD?
Figure 7: KL divergence comparison of gradient distributions against initial gradient distribution across training Steps. We use the projection weights in attention and MLP modules of the second layer as an example. The plots compare standard training with $\mathtt{GradNorm}$-augmented training. Lower KL divergence values indicate greater stability in gradient distributions.
To examine whether $\mathtt{GradNorm}$ stabilizes the distribution of the stochastic gradients as suggested by Theorem 1, we conduct controlled experiments using a scaled-down LLaMA-based model (about 10 million parameters) (Lialin et al., 2023), trained on the C4 dataset. Our goal is to measure how $\mathtt{GradNorm}$ affects the distribution of stochastic gradients over multiple training steps. Specifically, we employ a small-scale LLaMA-based model with approximately 10 million parameters  (Lialin et al., 2023). Training is conducted on the C4 dataset (Raffel et al., 2020).
Baselines
We compare:
  • Standard training: This uses an SGD optimizer with a learning rate of $5\times 10^{-4}$ and a linear learning rate scheduler, including a 10% warm-up of total training steps (10,000 steps).
  • $\mathtt{GradNorm}$-processed training: This applies $\mathtt{GradNorm}$ to pre-process the stochastic gradient before the parameter update. All other settings match the standard training baseline.
Methodology
We measure gradient statistics in the presence of mini-batch noise. At step $t=0$, we sample 16 additional mini-batches (batch size 64 each) and compute the mean and standard deviation of the corresponding raw or $\mathtt{GradNorm}$ gradients in each batch, and obtain the approximated initial gradient distribution. After each training step of baseline methods, we perform the same procedure and calculate the Kullback-Leibler divergence between the resulting gradient distributions and the initial gradient distributions. This process tracks how the gradient distribution changes over time.
Results
Figure 7 shows the KL divergence of gradient distributions for standard and $\mathtt{GradNorm}$-augmented training, relative to the corresponding initial approximated distributions. Apart from early spikes, $\mathtt{GradNorm}$ reduces fluctuations in the gradient distribution throughout training.
6.5.2 Does $\mathtt{GradWhitening}$ Counteracts Local Curvature and Provide Fast Convergence on ill-conditioned problems?
This subsection evaluates the optimization performance of gradient descent when combined with $\mathtt{GradWhitening}$. We use three classic problem settings:
(a) Quadratic optimization, well-conditioned
(b) Quadratic optimization, ill-conditioned
(c) Rastrigin function optimization
(d) Effect of initialization
  • High-dimensional quadratic optimization. A quadratic problem of the form in Equation 10, where ${\mathbf{W}}\in\mathbb{R}^{50\times 50}$.
  • Ill-conditioned quadratic optimization. Same setup as above, but with a deliberately chosen ill-conditioned ${\mathbf{H}}$.
  • Non-convex optimization with multiple local optima. We use the multivariate Rastrigin function:
    $f({\mathbf{W}})=m^{2}{\mathbf{A}}+\frac{1}{2}\text{Tr}[{\mathbf{W}}^{\top}{\mathbf{W}}]-{\mathbf{A}}\sum_{ij}\text{cos}(2\pi W_{ij}),$
    where ${\mathbf{W}}$ is an $m\times m$ matrix and $m=50$. This function has $10^{m^{2}}$ possible local optima.
Baselines
We compare five methods on all three problems: gradient descent (GD) with the theoretical optimal learning rate in Theorem 3, Adam with $\beta_{1}=0.9$, $\beta_{2}=0.999$ and a hand-tuned learning rate, Newton’s method with a tuned learning rate, and two $\mathtt{GradWhitening}$-based variants (with and without orthogonal initialization). This is to verify, under orthogonal initialization, $\mathtt{GradWhitening}$-processed GD behaves similarly to Newton’s method, as discussed in Section 5.3 and Theorem 2. All methods share the same initialization, except for the orthogonal $\mathtt{GradWhitening}$ variant, which projects the initial parameters onto an orthogonal matrix.
Results
From Figure 8 (a)–(c), we summarizes the following outcomes:
  • Quadratic problems (Figure 8 (a) and (b)). $\mathtt{GradWhitening}$ with orthogonal initialization and Newton’s method converge to optimum in one step, aligning with the theoretical predictions in Section 5.3 and Theorem 2.
  • Well- vs. ill-conditioned cases (Figure 8 (a) and (b)). In the well-conditioned setting (a), standard GD outperforms both Adam and $\mathtt{GradWhitening}$ (non-orthogonal initialization). In the ill-conditioned setting (b), $\mathtt{GradWhitening}$ (non-orthogonal initialization) outperforms GD by a large margin, while GD experiences slow convergence.
  • Comparison with Adam (Figure 8 (a)–(c)). In all three settings, $\mathtt{GradWhitening}$ with non-orthogonal initialization consistently outperforms Adam, consistent with Proposition 3.
  • Rastrigin function (Figure 8 (c)). On this non-convex problem, $\mathtt{GradWhitening}$ performs comparably to Newton’s method, with or without orthogonal initialization.
  • Effect of initialization (Figure 8 (d)) Furthermore, we force all methods to share the same orthogonalized initialization. As shown by the result, $\mathtt{GradWhitening}$ GD with orthogonal initialization still consistently outperforms all baselines, confirming that this initialization is only beneficial to $\mathtt{GradWhitening}$ GD among all baselines.